OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. m=8/9,(4,-3)

OpenStudy (anonymous):

Plug in the values into the equation: \[y = mx + b\] and find b. you will get the equation

OpenStudy (anonymous):

so it should look like this is that right. 4=8/9+-3 ?

OpenStudy (anonymous):

\[-3 = \frac{ 8 }{ 9 }(4) + b \] Solve for b.